Kevin invested part of his $10,000 bonus in a cd that paid 6% interest, and the rest in a fund that paid 11% interest. If his total interest was $900, how much did Kevin invest in the fund?
Answer by mananth(16949)

let hime invest $x in CD giving 6%
10000-x he invests in 11% fund
Total interest he received = $900
..
0.06x + 0.11(10,000-x) = 900
0.06x +1100-0.11x= 900
multiply by 100
6x+110000 -11x=90000
-5x=-20000
x= 4000.
Investment in 6% cd = $4000
$6000 at 11%